# Spend Wisely: Maximizing Post-Training Gains in Iterative Synthetic Data Boostrapping

### v1

@@ F.1 Toy Example
We show more results of the toy example with diverse parameters in Figure 5. All demonstrate the same conclusion as in Section 3.2.
@@ F.2 Ablation on Exponential Growth Schemes in Image Denoising
We conduct an ablation study for exponential growth policies to understand the impact of the $n_{t}$ on the performance of the iterative learning framework, as measured by PSNR in Figure 6. The results reveal that a smaller $n_{t}$ during the early stages of training leads to a rapid increase in PSNR and earlier convergence. In contrast, a larger $n_{t}$ results in a slower initial increase in PSNR but achieves a higher final PSNR when trained long enough. These findings underscore the trade-off between rapid early improvements and the potential for superior overall performance with prolonged training.
Figure 6: Ablation study for the image denoising task. We show the performance of exponential scheme under the $s=0.01$ setting with different parameters.
@@ F.3 Ablation on Exponential Growth Schemes in Math Reasoning
We further explore the effects of varying the parameter $n_{t}$ within exponential growth policies on the performance of mathematical reasoning tasks. As shown in Figure 7, configurations with higher $n_{t}$ demonstrate the potential for superior performance, improving the model’s reasoning capabilities and ability to solve mathematical problems. However, these configurations also exhibit some instability, characterized by abrupt changes in accuracy as $n_{t}$ increases. This highlights the need to find a balance between achieving rapid learning and mitigating the risk of unstable training dynamics. Based on these findings, we adopt $n_{t}=10\cdot 2^{t}\cdot B$ in the main text.
@@ F.4 Additional Empirical Results in Math Reasoning
In the main paper, we use a temperature of 0.3 for the math reasoning experiment shown in Figure 4. Here, we present additional empirical results with a temperature of 0.7 in Figure 8. The general trends across different curves and policy classes remain consistent when comparing these two figures. This ablation study further demonstrates the superiority of the proposed approach.
Figure 8: Additional empirical results of math reasoning where the temperature is 0.7. Other hyperparameters remain the same them in Figure 4, i.e., the training batch size is still set to $B=256$ and the exponential policy is still $n_{t}=10\cdot 2^{t}\cdot B$.

### v2

@@ F.1 Toy Example
Additional results for the toy example, obtained under various parameter settings, are shown in Figure˜5. In all cases, the behavior is consistent with the conclusion in Section˜3.2.
@@ F.2 Ablation on Exponential Growth Schemes in Image Denoising
We conduct an ablation study for exponential growth policies to understand the impact of the $n_{t}$ on the performance of the iterative learning framework, as measured by PSNR in Figure˜6. The results reveal that a smaller $n_{t}$ during the early stages of training leads to a rapid increase in PSNR and earlier convergence. In contrast, a larger $n_{t}$ results in a slower initial increase in PSNR but achieves a higher final PSNR when trained long enough. These findings underscore the trade-off between rapid early improvements and the potential for superior overall performance with prolonged training.
Figure 6: Ablation study for the image denoising task. We show the performance of exponential scheme under the $s=10$ setting with different parameters.
@@ F.3 Ablation on Exponential Growth Schemes in Math Reasoning
We further explore the effects of varying the parameter $n_{t}$ within exponential growth policies on the performance of mathematical reasoning tasks. As shown in Figure˜7, configurations with higher $n_{t}$ demonstrate the potential for superior performance, improving the model’s reasoning capabilities and ability to solve mathematical problems. However, these configurations also exhibit some instability, characterized by abrupt changes in accuracy as $n_{t}$ increases. This highlights the need to find a balance between achieving rapid learning and mitigating the risk of unstable training dynamics. Based on these findings, we adopt $n_{t}=10\cdot 2^{t}\cdot B$ in the main text.
@@ F.4 Additional Empirical Results in Math Reasoning
In the main paper, we use a temperature of 0.3 for the math reasoning experiment shown in Figure˜4. Here, we present additional empirical results with a temperature of 0.7 in Figure˜8. The general trends across different curves and policy classes remain consistent when comparing these two figures. This ablation study further demonstrates the superiority of the proposed approach.
Figure 8: Additional empirical results of math reasoning where the temperature is 0.7. Other hyperparameters remain the same them in Figure˜4, i.e., the training batch size is still set to $B=256$ and the exponential policy is still $n_{t}=10\cdot 2^{t}\cdot B$.
F.5 Mixed Policies
Table 3: Additional empirical results of mixed policies in the math reasoning task. For the mixed policies, we use a linear policy initially and switching to an exponential policy when the linear phase ceased to grow. The results show that mixed policies outperform linear policies and are comparable to the exponential policies.
Policy symbolic p1 p2
linear 64.24 40.84 19.17
mixed 65.98 45.56 20.16
exponential 65.66 47.26 20.65
We also explore mixed policies empirically. Specifically, we use a linear policy initially, switching to an exponential policy when the linear phase ceases to yield growth. We present the results in Table˜3, where we directly combine the linear and exponential schemes. The results show that mixed policies do not outperform the pure exponential policy. While this approach could be seen as a more general strategy, the increased degrees of freedom in parameter tuning make hyperparameter selection considerably more challenging. We leave further exploration of this direction to future work.
Appendix G Discussion
G.1 Impact Statement
This research advances the field of machine learning by addressing the strategic allocation of computational resources in the post-training phase using synthetic data. We introduce a novel theoretical framework to optimize the distribution of a fixed budget across multiple training iterations, an approach that demonstrates significant improvement in model performance. Our findings challenge conventional constant allocation strategies, revealing that dynamically adjusted policies, particularly those with exponential growth, not only converge more stably but also do so at a reduced computational cost. This work has broad implications for the development of AI models, especially in domains where the generation of synthetic data is critical for enhancing model capabilities without the extensive need for expensive or hard-to-obtain real-world data. The practical applications of this research extend to fields such as autonomous driving, medical image analysis, and any other area where AI must perform with high stability and efficiency under constraints of limited labeled data. This study thus provides a foundation for more sustainable AI development practices that prioritize both performance and computational efficiency.
G.2 Connection with Model Collapse
Prior work has introduced verification (Feng et al., 2024; El Firdoussi et al., 2025) and correction (Gillman et al., 2024) mechanisms for synthetic data to mitigate model collapse. They have theoretically derived one-step conditions under which collapse is avoided or triggered. Our methodology crucially relies on the reward model-based data verification process, using reward signals to filter data to prevent model collapse.
Empirically, we observe that increasing policies are generally more resistant to collapse. For constant policies, even with reward-based curation, the model’s performance initially improves but then degrades over successive iterations. Increasing polices yield smoother, generally upward trajectories, though some decline under very extensive finetuning. These observations extend beyond the scope of the analyses in (Feng et al., 2024; El Firdoussi et al., 2025) and suggest the need for a deeper theoretical analysis. We will add this discussion on model collapse to the paper and leave it for future work.
G.3 Limitation
In this paper, we focus on achieving efficient SFT improvement within an iterative learning framework. Recent trends (Guo et al., 2025) incorporate reinforcement learning into the entire post-training phase, while SFT remains an important component for enabling the model to exhibit certain behaviors and skills necessary for successful RL training later (Gandhi et al., 2025). We aim to extend the theory to develop an efficient training scheme for RLHF and RL from verifiable rewards.
We also examine learning with access to a reward model with (Section˜B.5) or without independent random noises. An empirical reward, such as using an LLM as a judge, could produce biased results and lead to reward hacking. We leave the exploration of biased rewards to future work.
We conduct two experiments to validate our theoretical results. However, incorporating the SFT scheme design into production scale is beyond our resources at academic institutions. We present theoretical results and experimental findings across modalities that support the generalizability of our results.